CAVITY RINDDOWN SPECTROSCOPY OF THE Ã−X̃ ELECTRONIC TRANSITION OF THE CH_{3}C(O)O_{2} RADICAL

\underline{SERGEY J. ZALYUBOVSKY}, BRENT G. GLOVER, TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus OH 43210.

The acetyl peroxy radical, CH_{3}C(O)O_{2}, is one of the most abundant organic peroxy radicals in the atmosphere ^{a}. It plays a significant role in many key atmospheric processes, including the atmospheric ozone balance ^{b} and the formation of peroxy acetyl nitrate ^{c}, a key component in photochemically generated smog. Various advantages in spectroscopic and kinetic applications can be gained by studying peroxy radicals via their near IR electronic transition ^{d}^{,}^{e}. In this talk we report cavity ringdown spectra of the Ã^{2}A′ - X̃^{2}A′′ electronic transition of acetyl peroxy radical and its perdeutero analog. For CH_{3}C(O)O_{2}, a T_{00} electronic origin is observed at 5582.5(5) cm^{−1}, and several Ã state vibrational frequencies are measured and assigned. Extensive ab initio calculations have been carried out to confirm assignments for the electronic origin and observed vibrational hot bands. The empirical absorption cross section for the origin of Ã^{2}A′ - X̃^{2}A′′ transition has been estimated.

aG. S. Tyndall, R. A. Cox, C. Granier, R. Lesclaux, G. K. Moortgat, M. J. Pilling, A. R. Ravishankara and T. J. Wallington, J. Geophys. Research [Atmospheres] 106(D11), 12157(2001)

bP. D. Lightfoot, R. A. Cox, J. N. Crowley, M. Destriau, G. D. Hayman, M. E. Jenkin, G. K. Moortgat and F. Zabel, Atmos. Environ. A-Gen. 26A, 1805(1992)

cH. J. Beine, D. A. Jaffe, J. A. Herring, J. A. Kelley, T. Krognes and F. Stord, J. Atmos. Chem. 27, 127(1997)

dM. B. Pushkarsky, S. J. Zalyubovsky and T. A. Miller, J. Chem. Phys. 112, 10695(2000)

eS. J. Zalyubovsky, Dongbing Wang and T. A. Miller, Chem. Phys. Lett. 335, 298(2001)